### v2

# Emergent Weyl-like points in periodically modulated systems and Floquet engineering

Abstract
We investigate a three-dimensional topological phase resembling a Weyl semimetal, modulated by a periodic potential and engineered through Floquet dynamics. This system is constructed by stacking two-dimensional Chern insulators and hosts Weyl-like points defined in the parameter space $(k_{x},k_{y},z)$, distinct from conventional Weyl points in momentum space $(k_{x},k_{y},k_{z})$. The Weyl-semimetal-like phase exhibits characteristics akin to those of Weyl semimetals, including linear dispersion near the Weyl-like points, nontrivial bulk topology, and the presence of Fermi arcs connecting the Weyl-like points. Unlike traditional Weyl semimetals, these features manifest in real space rather than momentum space. Additionally, we compute the layer-resolved Floquet Hall conductivity and demonstrate that the positions of the Weyl-like points can be controlled via high-frequency laser pumping.
I Introduction
Stacks of two-dimensional (2D) Chern insulator phases have proven to be a versatile platform for realizing extraordinary topological phases, attracting considerable attention in recent years Mogi et al. (2017a, b); Xiao et al. (2018); Mogi et al. (2022); Qin et al. (2023); Chen and Shen (2024); Chang et al. (2013); Yoshimi et al. (2015); Shan et al. (2010); Lu et al. (2010); Li et al. (2010); Lu et al. (2013); Chen et al. (2019); Sun et al. (2020); Liu et al. (2010); Dabiri et al. (2021); Dabiri and Cheraghchi (2021); Sun et al. (2023); Qin et al. (2022a); Zhu et al. (2023). For instance, the well-known Weyl semimetal phase can be interpreted as a stack of 2D Chern insulators with orbital-dependent interlayer tunneling Yang et al. (2011); Chen et al. (2015, 2018a, 2018b); Li et al. (2021). Similarly, phases such as the axion insulator and the hinged quantum spin Hall effect arise from alternating stacks of Chern insulators with opposite Chern numbers Chen et al. (2021, 2024, 2023a, 2023b). Moreover, higher-Chern-number quantum anomalous Hall insulators can be constructed by alternating stacks of Chern insulators and normal insulators Zhao et al. (2022); Baba et al. (2022); Li et al. (2022, 2023).
Weyl semimetals are characterized by linearly dispersing band-touching points, known as Weyl points, in momentum space Wan et al. (2011); Chan et al. (2016); Yan and Felser (2017); Guo et al. (2023). These Weyl points are topologically protected and act as monopoles or antimonopoles of the Berry curvature Li et al. (2016); Lu and Shen (2017). According to the no-go theorem Nielsen and Ninomiya (1981a, b), Weyl points always appear in pairs and are connected by surface states known as Fermi arcs.
Recent advancements have realized ideal Weyl semimetals across various platforms. They have been implemented in electrical circuits Lu et al. (2019); Lee et al. (2020), ultracold atomic gases with three-dimensional (3D) spin-orbit coupling Dubček et al. (2015); Roy et al. (2018); Wang et al. (2021); Lu et al. (2020), photonic crystals Lu et al. (2015); Li et al. (2018), and acoustic metamaterials Xiao et al. (2015); Yang and Zhang (2016); Peri et al. (2019).
Zhang et al. recently proposed and studied a 2D phase modulated by a periodic potential Zhang et al. (2023). This system exhibits edge modes with spectral flow dependent on position, in contrast to traditional Chern insulators, where the spectral flow of edge modes depends on momentum. Moreover, they suggested that the periodic potential could be experimentally realized through the effects of charge density waves Huang et al. (2021); Qin et al. (2020).
Off-resonant light serves as a powerful tool for inducing Floquet Weyl semimetals Wang et al. (2014); Hübener et al. (2017). Building on Floquet theory Oka and Aoki (2009); Zhan et al. (2024); Qin et al. (2023, 2022b, 2022a, 2024); Dabiri et al. (2021); Dabiri and Cheraghchi (2021); Dabiri et al. (2022); Cheraghchi and Askarpour (2023); Pervishko et al. (2018); Calvo et al. (2015); Seshadri and Sen (2022); Seshadri et al. (2019); Seshadri (2023); Lee et al. (2018); Magnus (1954); Blanes et al. (2009); Bukov et al. (2015); Eckardt and Anisimovas (2015); Chen et al. (2018c, d); Du et al. (2022); Wang et al. (2023), it would be intriguing to investigate the experimental implications of realistic periodic driving frequencies. This is particularly relevant in light of recent developments, such as the photoinduced anomalous Hall effect in graphene McIver et al. (2020) and its corresponding microscopic theory Sato et al. (2019), the experimental observation of Floquet states in graphene Merboldt et al. (2024), and the direct detection of Floquet-Bloch states in monolayer graphene Choi et al. (2024), WSe2 Aeschlimann et al. (2021), and the 3D topological insulator Bi2Se3 Wang et al. (2013). Additionally, experimental studies on Floquet transport effects, including anomalous linear transport in light-driven 3D Dirac electrons in bismuth Hirai et al. (2023) and nonperturbative nonlinear transport in a magnetic Floquet-Weyl semimetal Day et al. (2024), further highlight the potential of this approach.
In this work, we propose and study a 3D topological phase resembling a Weyl semimetal, modulated by a periodic potential and realized through a simple stack of 2D Chern insulators. The system features Weyl-like points defined in the parameter space $(k_{x},k_{y},z)$, distinct from the conventional Weyl points in momentum space $(k_{x},k_{y},k_{z})$. This Weyl-semimetal-like phase exhibits key properties of a Weyl semimetal, including linear dispersion near the Weyl-like points, nontrivial bulk topology, and the emergence of Fermi arcs connecting these points. However, these behaviors are observed in real space, rather than momentum space, as in traditional Weyl semimetals. Additionally, we calculate the layer-resolved Floquet Hall conductivity under high-frequency periodic driving and show that the positions of the Weyl-semimetal-like nodes can be tuned by high-frequency laser driving.
The structure of this paper is as follows: In Section II, we introduce the Weyl-like lattice Hamiltonian with a periodic potential along the $z$-direction, along with the conventional Weyl semimetal model, and discuss the duality between these two models. In Section III, we derive the Floquet effective Hamiltonians for both models under high-frequency laser pumping. Section IV presents calculations of the Floquet Weyl nodes, the layer Floquet Chern number, the Floquet energy spectrum, and the Floquet Fermi arc for both models under high-frequency laser driving. A brief discussion on potential experimental realizations of the Weyl-like model is provided in Section V. Finally, Section VI offers a summary of the key results.
II Model
We begin with the two-band model Hamiltonian describing a stack of 2D Chern insulators in a periodic potential along the $z$-direction,
$H\!=\!\sum_{j_{z}}\left[h_{\text{2D}}({\bf k}_{||})\!+\!V(z)\sigma_{z}\right]c_{j_{z}}^{{\dagger}}c_{j_{z}},$ (1)
where $c_{j_{z}}^{{\dagger}}$ is the electron creation operator in the $j_{z}$-th layer, $h_{\text{2D}}({\bf k}_{||})\!=\!m_{z}\sigma_{z}\!+\!m_{0}[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)]\sigma_{z}\!+\!t_{x}\sin(k_{x}a)\sigma_{x}\!+\!t_{y}\sin(k_{y}a)\sigma_{y}$ represents the Chern insulator Hamiltonian when $m_{z}\times m_{0}<0$, and ${\bf k}_{||}\!=\!(k_{x},k_{y})$. Here, $a$ is the lattice constant, and $\sigma_{i}$ $(i\!=\!x,y,z)$ are Pauli matrices acting on the orbital space. The spatially periodic potential is given by $V(z)\!=\!V_{z}\cos(Q_{z}z)$, where $V_{z}$ is the potential strength, $z=j_{z}a$ is the position along the $z$-direction, and $\lambda_{z}\!=\!2\pi/Q_{z}$ is the periodic wavelength. In the following calculations, we use the fixed parameters $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!t_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, and $\lambda_{z}\!=\!31a$ with $a\!=\!1$ nm. The detailed derivation of the matrix form for the real-space model in a periodic potential along the $z$-direction, with open boundary conditions (OBCs) in $y$-direction and periodic boundary conditions (PBCs) in $x$-direction, is provided in Section SI in the Supplemental Material Sup .
Figure 1: Duality between the decoupled Weyl-like model (left column) and the Weyl semimetal model (right column). (a1) Schematic representation of the decoupled Weyl-like model in real space, featuring a spatially periodic potential $V(z)=V_{z}\cos(Q_{z}z)$, where $Q_{z}=2\pi/\lambda_{z}$. (b1) Weyl-like cones in $(k_{x},k_{y},z)$ space at two critical planes, $(k_{x},k_{y},z_{c-})$ and $(k_{x},k_{y},z_{c+})$, marking topological transitions. (c1) Layer-resolved Chern number as a function of layer $z$ for the decoupled Weyl-like model with $V_{z}=1$ eV, $\lambda_{z}=L_{z}=31a$, and the lattice constant $a=1$ nm. (d1) Energy spectrum (blue solid curves) for the Weyl-like model with $k_{x}=k_{y}=0$. (e1) Energy spectrum (blue solid curves) and Fermi-like arc (blue horizontal line for $E=0$) for the Weyl-like model under OBCs along the $y$-direction, with $k_{x}=0$ and system size $L_{y}=31a$. (a2) Schematic representation of the Weyl semimetal model in real space, with a momentum-periodic potential $U(k_{z})=t_{z}\cos(k_{z}a)$. (b2) Weyl cones in the $(k_{x},k_{y},k_{z})$ momentum space at two critical planes $(k_{x},k_{y},-k_{c})$ and $(k_{x},k_{y},+k_{c})$, indicating the topological features of the model. (c2) Chern number as a function of momentum $k_{z}$ for the Weyl semimetal model with $t_{z}=1$ eV. (d2) Energy spectrum (black solid curves) for the Weyl semimetal model with $k_{x}=k_{y}=0$. (e2) Energy spectrum (black solid curves) and Fermi arc (black horizontal line for $E=0$) for the Weyl semimetal model under OBCs along the $y$-direction, with $k_{x}=0$ and system size $L_{y}=31a$. The other parameters are $m_{z}=0.5$ eV, $m_{0}=1$ eV, $t_{x}=t_{y}=t_{0}=1$ eV, and $L_{z}=31a$.
II.1 Weyl-like model
The spectral flow of the system is described by the Weyl-like model Hamiltonian density, Eq. (1), defined in the $(k_{x},k_{y},z)$-space as follows:
[…]
$\!+t_{x}\sin(k_{x}a)\sigma_{x}\!+\!t_{y}\sin(k_{y}a)\sigma_{y}.$
As shown in Fig. 1(a1), the system can be interpreted as an array of isolated 2D systems, each with a different topological mass: $\tilde{m}_{z}(z)\!=\!m_{z}\!+\!V(z)$, where $V(z)\!=\!V_{z}\cos(Q_{z}z)$.
To study the topological properties of the system, we define the layer-resolved Chern number for the $m$-th occupied band as Fukui et al. (2005); Sun et al. (2020); Otrokov et al. (2019); Gao et al. (2021); Chen et al. (2024); Mong et al. (2010); Essin et al. (2009); Wang et al. (2015); Varnava and Vanderbilt (2018); Chen and Shen (2024); Fu et al. (2021); Price et al. (2015); Zhang et al. (2019)
$C_{m}(z)\!=\!\frac{1}{2\pi}\int{\cal F}_{xy}^{mm}({\bf k}_{||},z)d^{2}{\bf k}_{||},$ (3)
where the non-Abelian Berry curvature is given by Chen and Shen (2024); Fu et al. (2021); Price et al. (2015); Zhang et al. (2019)
[…]
$\!+i[{\cal A}_{k_{x}}^{mn}({\bf k}_{||},z),{\cal A}_{k_{y}}^{mn}({\bf k}_{||},z)],$
where ${\cal A}_{k_{\alpha}}^{mn}({\bf k}_{||},z)$ is the Berry connection of the occupied bands with $m,n\in{\rm occ}$, defined as
${\cal A}_{k_{\alpha}}^{mn}({\bf k}_{||},z)\!=\!i\langle\psi_{m}|\partial_{k_{\alpha}}\psi_{n}\rangle.$ (5)
Here, “occ” is used to denote the occupied bands. A detailed derivation of the non-Abelian Berry curvature in (4) is provided in Section SII in the Supplemental Material Sup .
The layer-resolved Chern number is illustrated in Fig. 1(c1). Figures 1(a1) and 1(b1) show that each 2D system behaves as a Chern insulator with a Chern number $|C|=1$ for $z_{c-}<z<z_{c+}$, and as a normal insulator with $C=0$ for $z<z_{c-}$ and $z>z_{c+}$. The critical values of $z$, denoted by $z_{c\pm}$, are determined by the condition $\tilde{m}_{z}(z)\!=\!m_{z}\!+\!V(z)\!=\!0$ and are given by:
$\frac{z_{c\pm}}{\lambda_{z}}\!=\!\frac{1}{2\pi}\left[\pi\pm\arccos\left(\frac{m_{z}}{V_{z}}\right)\right],$ (6)
where the condition $-1\leqslant\tilde{m}_{z}/V_{z}\leqslant 1$ is satisfied.
Figure 1(d1) shows the energy spectrum as a function of position $z$ with $k_{x}\!=\!k_{y}\!=\!0$. Notably, a band inversion is observed near $z\!=\!z_{c\pm}$, clearly illustrating the spectral flow of the stacked system. Additionally, the edge states (zero-energy modes) can be observed in Fig. 1(e1) that are dual to Fig. 1(e2).
II.2 Weyl semimetal
To facilitate a comparative study, we review the Weyl semimetal model Hamiltonian, which takes the following form Yang et al. (2011); Chen et al. (2015, 2018a, 2018b)
[…]
$\!-m_{0}\left[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)\right]\sigma_{z}$
$\!+t_{x}\sin(k_{x}a)\sigma_{x}\!+\!t_{y}\sin(k_{y}a)\sigma_{y},$
where the Hamiltonian describes the Weyl semimetal phase, which can be interpreted as a stack of 2D materials with orbital-dependent hopping $t_{z}$, as illustrated in Fig. 1(a2). The matrix form of this Hamiltonian (7) in real space, with OBCs in the $z$-direction and PBCs in the $x$- and $y$-directions, is provided in Subsection SIII A in the Supplemental Material Sup . Similarly, the matrix form with OBCs in the $y$-direction and PBCs in the $x$- and $z$-directions is given in Subsection SIII B in the Supplemental Material Sup .
Rewriting Eq. (7) in the form ${\cal H}_{\rm W}({\bf k})\!=\!\mathbf{h}({\bf k})\cdot{\boldsymbol{\sigma}}$, where $\mathbf{h}({\bf k})\!=\!(h_{x},h_{y},h_{z})$, we define the components as: $h_{x}\!=\!t_{x}\sin(k_{x}a)$, $h_{y}\!=\!t_{y}\sin(k_{y}a)$, $h_{z}\!=\!M_{z}(k_{z})\!-\!m_{0}\left[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)\right]$, with $M_{z}(k_{z})\!=\!-m_{z}\!+\!t_{z}\cos(k_{z}a)$. The eigenenergies of this 2D model Hamiltonian (7) are given by
$E_{{\rm W},\pm}\!=\!\pm\sqrt{h_{x}^{2}\!+\!h_{y}^{2}+h_{z}^{2}}.$ (8)
Thus, the energy gap between the top and bottom surfaces at $k_{x}\!=\!k_{y}\!=\!0$ is: $\Delta_{\rm W}\!=\!E_{{\rm W},+}\!-\!E_{{\rm W},-}\!=\!2|M_{z}(k_{z})|$ as shown in Fig. 1(d2).
The Chern number of the Weyl semimetal model (7) can then be calculated using the expression Thouless et al. (1982); Fukui et al. (2005); Xiong et al. (2016); Liu et al. (2019); Lu et al. (2010); Shan et al. (2010); Qin et al. (2022a)
$C(k_{z})\!=\!\frac{1}{4\pi}\int_{\text{B.Z.}}\!\!d^{2}{\bf k}_{||}\frac{\mathbf{h}({\bf k})}{|\mathbf{h}({\bf k})|^{3}}\!\cdot\![\partial_{k_{x}}\mathbf{h}({\bf k})\!\times\!\partial_{k_{y}}\mathbf{h}({\bf k})],$ (9)
where ${\bf k}_{||}\!=\!(k_{x},k_{y})$ and the integral is over the 2D Brillouin zone. A detailed analytical expression for the Chern number (9) of the Weyl semimetal model (7) is provided in Section SIV in the Supplemental Material Sup .
For the special case where $k_{x}\!=\!k_{y}\!=\!0$, the eigenenergies of the Weyl semimetal Hamiltonian (7) are:
$E_{\pm}\!=\!\pm\left[-m_{z}\!+\!t_{z}\cos(k_{z}a)\right].$ (10)
In this case, the system hosts a pair of gapless Weyl nodes at $(0,0,k_{c\pm})$ with the positions $k_{c\pm}\!=\!\pm\frac{1}{a}\arccos(m_{z}/t_{z})$, under the condition $|m_{z}/t_{z}|\leqslant 1$ as shown in Figs. 1(b2) and 1(c2). These Weyl nodes $k_{c}$ are obtained by solving $M_{z}(k_{z})\!=\!-m_{z}\!+\!t_{z}\cos(k_{z}a)\!=\!0$ with the condition $-1\leqslant m_{z}/t_{z}\leqslant 1$. For $k_{z}$ values within the range $k_{c-}<k_{z}<k_{c+}$, the system behaves as a Chern insulator with Chern number $C\!=\!1$. Outside this range, for $k_{z}<k_{c-}$ or $k_{z}>k_{c+}$, the system becomes a normal insulator with $C\!=\!0$, as depicted in Fig. 1(c2).
Figure 2: (a) Weyl-like nodes at $z\!=\!z_{c-}$ and $z\!=\!z_{c+}$, marking topological transitions. (b) The two Weyl-like nodes are separated by the bulk gap between $z\!=\!z_{c-}$ and $z\!=\!z_{c+}$. Each half of the system contains a single Weyl-like node. This differs from conventional Weyl points, which typically appear in pairs in momentum space $(k_{x},k_{y},k_{z})$.
II.3 Comparison between the two models
As shown in Fig. 1, the two models – the decoupled Weyl-like model (left column) and the Weyl semimetal model (right column) – can be viewed as duals of each other. The Weyl-like model in Eq. (2) provides a generic framework for describing Weyl semimetals, including key features such as band touching, Chern number, and Fermi arcs, all of which are captured by the Weyl semimetal model in Eq. (7). Figure 2 illustrates that the two Weyl-like nodes are separated by the bulk gap between $z\!=\!z_{c-}$ and $z\!=\!z_{c+}$, with each half of the system hosting a single Weyl-like node. This contrasts with conventional Weyl points, which, due to the no-go theorem Nielsen and Ninomiya (1981a, b), typically appear in pairs in momentum space with opposite chiralities. Below, we will focus on the case involving Floquet engineering.
III Floquet Hamiltonian
To investigate the Floquet engineering of the two models discussed in Section II, we derive their effective Floquet Hamiltonians.
The optical field propagating along the $z$-direction in the material is given by ${\bf E}(t)=\partial{\bf A}(t)/\partial t=E_{0}(\cos(\omega t),\cos(\omega t+\varphi),0)$, where $E_{0}$ is the amplitude, $\omega$ is the angular frequency, and $\varphi$ represents the phase. For linear polarization, $\varphi=0$, while $\varphi=\mp\pi/2$ corresponds to left- or right-handed circular polarization. The vector potential is periodic with period $T=2\pi/\omega$, and it is given by ${\bf A}(t)={\bf A}(t+T)=\omega^{-1}E_{0}(\sin(\omega t),\sin(\omega t+\varphi),0)$. In the off-resonant regime, where the central Floquet band is well-separated from other replicas, the high-frequency expansion applies. We choose the optical frequency as $\hbar\omega\!=\!10$ eV, which is much larger than the material’s bandwidth Zhan et al. (2024); Qin et al. (2023, 2022b, 2022a, 2024); Dabiri et al. (2021); Dabiri and Cheraghchi (2021); Dabiri et al. (2022); Cheraghchi and Askarpour (2023); Pervishko et al. (2018).
Under optical driving, the electron motion is modified by minimal coupling to the electromagnetic gauge field ${\bf A}(t)$, resulting in the photon-dressed effective Hamiltonian:
${\cal H}({\bf k},t)\!=\!{\cal H}\left({\bf k}\!-\!\frac{e}{\hbar}{\bf A}(t)\right).$ (11)
Next, we derive the effective Floquet Hamiltonian Oka and Aoki (2009); Calvo et al. (2015); Seshadri and Sen (2022); Seshadri et al. (2019); Seshadri (2023); Lee et al. (2018)
[…]
where $\mathcal{T}$ is the time-ordering operator. In the high-frequency regime, a closed-form solution exists via the high-frequency expansion Magnus (1954); Blanes et al. (2009); Lee et al. (2018); Bukov et al. (2015); Eckardt and Anisimovas (2015); Chen et al. (2018c, d); Du et al. (2022); Wang et al. (2023):
${\cal H}^{(F)}({\bf k})\!=\!{\cal H}_{0}\!+\!\sum_{n=1}^{\infty}\frac{[{\cal H}_{-n},{\cal H}_{n}]}{n\hbar\omega}\!+\!{\cal O}(\omega^{-2}),$ (13)
[…]
III.1 Floquet Weyl-like model
From Eq. (13), the Floquet Hamiltonian for the Weyl-like model can be expressed as
${\cal H}^{(F)}({\bf k}_{||},z)\!=\!\sum_{i=x,y,z}h_{i}\sigma_{i},$ (14)
where
[…]
$\!-\!\!\!\!\!\!\!\sum_{n\in{\rm odd},n>0}\!\frac{4t_{x}m_{0}{\cal J}_{n}^{2}(A_{0}a)}{n\hbar\omega}\sin(n\varphi)\sin(k_{x}a)\cos(k_{y}a),$
[…]
$\!-\!\!\!\!\!\!\!\sum_{n\in{\rm odd},n>0}\!\frac{4t_{x}m_{0}{\cal J}_{n}^{2}(A_{0}a)}{n\hbar\omega}\sin(n\varphi)\cos(k_{x}a)\sin(k_{y}a),$
[…]
$\!+{\cal J}_{0}(A_{0}a)m_{0}\left[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)\right]$
$\!+\!\!\!\!\!\!\!\sum_{n\in{\rm odd},n>0}\!\frac{4t_{x}t_{y}{\cal J}_{n}^{2}(A_{0}a)}{n\hbar\omega}\sin(n\varphi)\cos(k_{x}a)\cos(k_{y}a),$
${\cal J}_{n}(A_{0}a)$ is the $n$th Bessel function of the first kind Temme (1996), and $A_{0}\!=\!eE_{0}/(\hbar\omega)$. The detailed derivations leading to Eq. (14) are provided in Subsection SV A in the Supplemental Material Sup . For $\varphi\!=\!0$ (linearly polarized laser), the Weyl-like nodes are determined by solving the equation $m_{z}\!+\!V_{z}\cos(Q_{z}z)\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]\!=\!0$, subject to the condition $-1\leqslant\cos(Q_{z}z)\!=\!\{m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]\}/V_{z}\leqslant 1$. This yields the Weyl-like nodes at
$\frac{z_{c\pm}}{\lambda_{z}}\!=\!\frac{1}{2\pi}[\pi\pm\arccos(\tilde{m}_{z}(A_{0})/V_{z})],$ (18)
where $\tilde{m}_{z}(A_{0})\!=\!m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]$ and $-1\leqslant\tilde{m}_{z}(A_{0})/V_{z}\leqslant 1$. By substituting the model parameters $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, and $a\!=\!1$ nm into the condition $|\tilde{m}_{z}(A_{0})/V_{z}|<1$, we have $A_{0}<1.034$ nm-1. From Eq. (18), it is clear that the position of the Weyl-like nodes can be tuned by adjusting the intensity of the optical field.
Additionally, Section SVI of the Supplemental Material Sup presents the derivations for the Weyl-like model under PBCs along the $x$-direction and OBCs along the $y$-direction.
III.2 Floquet Weyl semimetal
The Floquet Hamiltonian for the Weyl semimetal, derived from Eq. (13), takes the form:
${\cal H}^{(F)}_{\rm W}({\bf k})\!=\!\sum_{i=x,y,z}h_{i}\sigma_{i},$ (19)
where
[…]
$\!+\!\!\!\!\!\!\!\sum_{n\in{\rm odd},n>0}\!\frac{4t_{x}m_{0}{\cal J}_{n}^{2}(A_{0}a)}{n\hbar\omega}\sin(n\varphi)\sin(k_{x}a)\cos(k_{y}a),$
[…]
$\!+\!\!\!\!\!\!\!\sum_{n\in{\rm odd},n>0}\!\frac{4t_{x}m_{0}{\cal J}_{n}^{2}(A_{0}a)}{n\hbar\omega}\sin(n\varphi)\cos(k_{x}a)\sin(k_{y}a),$
[…]
$\!-{\cal J}_{0}(A_{0}a)m_{0}\left[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)\right]$
$\!+\!\!\!\!\!\!\!\sum_{n\in{\rm odd},n>0}\!\frac{4t_{x}t_{y}{\cal J}_{n}^{2}(A_{0}a)}{n\hbar\omega}\sin(n\varphi)\cos(k_{x}a)\cos(k_{y}a).$
A detailed derivation of Eq. (19) is provided in Subsection SV B in the Supplemental Material Sup . For linearly polarized light ($\varphi\!=\!0$), the Weyl nodes are determined by solving the equation $-m_{z}\!+\!t_{z}\cos(k_{z}a)\!-\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]\!=\!0$, subject to the condition $-1\leqslant\cos(k_{z}a)\!=\!\{m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]\}/t_{z}\leqslant 1$. This condition gives the Weyl node positions:
$k_{c\pm}\!=\!\pm\frac{1}{a}\arccos(\tilde{m}_{z}(A_{0})/t_{z}),$ (23)
where $\tilde{m}_{z}(A_{0})\!=\!m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]$ and $-1\leqslant\tilde{m}_{z}(A_{0})/t_{z}\leqslant 1$. Substituting typical model parameters, $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{z}\!=\!1$ eV, and $a\!=\!1$ nm, into the condition $|\tilde{m}_{z}(A_{0})/t_{z}|<1$, we obtain the constraint $A_{0}<1.034$ nm-1. From Eq. (23), we can conclude that the position of the Weyl nodes can be tuned by adjusting the intensity of the optical field.
The detailed derivation of the Floquet Hamiltonian matrix for the Weyl semimetal, under OBCs along the $z$-direction and PBCs along the $x$- and $y$-directions, is provided in Subsection SVII A of the Supplemental Material Sup . Similarly, Subsection SVII B of the Supplemental Material Sup presents the derivation for the case of PBCs along the $x$- and $z$-directions and OBCs along the $y$-direction.
III.3 Validity
[…]
IV Numerical results for Floquet engineering
In this section, we present the numerical results for Floquet engineering, with a comparative analysis between the Floquet Weyl-like and Floquet Weyl semimetal models.
IV.1 Floquet Weyl nodes
Figure 3: (a) Floquet Weyl-like nodes and (b) Floquet Weyl nodes as a function of light intensity $A_{0}$. The parameters are $\varphi=0$ (linearly polarized laser), $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!1$ eV, $V_{z}\!=\!1$ eV, $t_{z}\!=\!1$ eV, and $L_{z}\!=\!\lambda_{z}\!=\!31a$.
As shown in Fig. 3, the positions of the Floquet Weyl-like nodes [Fig. 3(a)] and the Floquet Weyl nodes [Fig. 3(b)] can be controlled by the light intensity $A_{0}$. Particularly, when $A_{0}$ reaches a threshold value of approximately $\sim 1.034$ nm-1, the two Floquet Weyl-like nodes (or Floquet Weyl nodes) merge into a single node, signaling a topological transition.
IV.2 Floquet Chern number
Figure 4: Chern numbers for (a1,b1) the Floquet Weyl-like model and (a2,b2) the Floquet Weyl semimetal. (a1,a2) $A_{0}\!=\!0.1$ nm-1; (b1,b2) $A_{0}\!=\!0.9$ nm-1. The other parameters are the same as those in Fig. 3.
We compute the Chern numbers for both the Floquet Weyl-like model and the Floquet Weyl semimetal model, as shown in Fig. 4. The results indicate that the phase transition points, represented by the Floquet Weyl nodes, shift with varying light intensities. Detailed derivations of the Chern numbers for both Floquet models are provided in Sections SVIII and SIX in the Supplemental Material Sup , respectively.
IV.3 Floquet energy spectrum
Figure 5: Energy spectra for (a1,b1) the Floquet Weyl-like model and (a2,b2) the Floquet Weyl semimetal with $k_{x}\!=\!k_{y}\!=\!0$. (a1,a2) $A_{0}\!=\!0.1$ nm-1; (b1,b2) $A_{0}\!=\!0.9$ nm-1. The other parameters are the same as those in Fig. 3.
The energy spectra for the Floquet Weyl-like model and the Floquet Weyl semimetal, calculated at $k_{x}=k_{y}=0$, are shown in Fig. 5. The energy gapless points, corresponding to the Weyl nodes, exhibit a dependence on the light intensity.
IV.4 Floquet Fermi arc
Figure 6: Energy spectra (solid curves) and Fermi arcs (horizontal straight line for $E\!=\!0$) for (a1,b1) the Floquet Weyl-like model and (a2,b2) the Floquet Weyl semimetal under OBCs along the $y$ direction with $k_{x}\!=\!0$ and $L_{y}\!=\!31a$. (a1,a2) $A_{0}\!=\!0.1$ nm-1; (b1,b2) $A_{0}\!=\!0.9$ nm-1. The other parameters are the same as those in Fig. 3.
In Fig. 6, we show the energy spectra (solid curves) and Fermi arcs (depicted as horizontal straight lines at $E=0$) for both the Floquet Weyl-like model and the Floquet Weyl semimetal under OBCs along the $y$-direction, with $k_{x}=0$ and $L_{y}=31a$. We observe that the lengths of the Floquet Fermi arcs change with light intensity, reflecting variations in the Floquet Weyl nodes.
V Experimental Realizations
In this section, we discuss potential experimental realizations of the Weyl-like model. Weyl semimetals have been successfully realized in various platforms, including electrical circuits Lu et al. (2019); Lee et al. (2020), ultracold atomic gases Wang et al. (2021); Lu et al. (2020), photonic crystals Lu et al. (2015); Li et al. (2018), and acoustic metamaterials Xiao et al. (2015); Yang and Zhang (2016); Peri et al. (2019). For the purpose of illustration, we focus on the realization of the Weyl-like model using electrical circuits.
Unlike the 3D cubic lattices employed in cold-atom proposals Dubček et al. (2015); Roy et al. (2018) and experimental realizations Wang et al. (2021); Lu et al. (2020), we propose a design based on a stack of 2D square lattices, with each plaquette in every principal plane hosting a flux of $\pi$, as described in Ref. Ningyuan et al. (2015). This configuration corresponds to a system of spin-$1/2$ particles with engineered spin-dependent tunneling.
The system thus takes the form of a multi-layer 2D time-reversal symmetric tight-binding model, with two sites per unit cell. This differs from the 3D time-reversal symmetric tight-binding model on a cubic lattice, which was originally proposed in Ref. Lu et al. (2019) to realize Weyl semimetals in 3D momentum space.
To implement this multi-layer 2D square lattice in an electrical circuit, each lattice site is represented by an inductor, while tunnel couplings are modeled by pairs of capacitors. Zero-phase tunnel couplings connect the positive ends of adjacent inductors, while their negative ends are similarly linked. A $\pi$-phase tunnel coupling capacitively links the positive end of one inductor to the negative end of a neighboring inductor, and vice versa.
The stacking of the 2D square lattice layers is coupled with a spatially periodic potential, which can be induced by an alternating current Qin et al. (2024) or via the charge-density-wave mechanism Zhang et al. (2023); Qin et al. (2020). This approach offers a promising platform for simulating the Weyl-like model, enabling experimental exploration of its novel topological properties.
VI Summary
In our study, we introduce a novel 3D topological phase with Weyl-semimetal-like characteristics, which is induced by a periodic potential. This phase is realized by stacking 2D Chern insulators, leading to Weyl-like points in the parameter space $(k_{x},k_{y},z)$, in contrast to the conventional Weyl points in momentum space $(k_{x},k_{y},k_{z})$.
The Weyl-semimetal-like phase we propose shares key features with traditional Weyl semimetals, such as linear dispersion near the Weyl-like points, nontrivial bulk topology, and the emergence of Fermi arcs that connect these points. Importantly, these topological behaviors manifest in real space, rather than in momentum space.
Additionally, our investigation explores the effects of high-frequency periodic driving, where we calculate Floquet Weyl nodes, layer Floquet Chern numbers, Floquet energy spectra, and Floquet Fermi arcs. We demonstrate that the Weyl-semimetal-like nodes can be tuned via high-frequency laser driving, providing a means to manipulate these topological features. This work opens new possibilities for exploring and controlling topological phases through external stimuli.
Acknowledgements.
We acknowledge helpful discussions with Ching Hua Lee. R.C. acknowledges the support from the National Natural Science Foundation of China (Grant No. 12304195) and the Chutian Scholars Program in Hubei Province.